# ROLCH: Regularized Online Learning for Conditional Heteroskedasticity

### v1

@@ 6 Conclusion
We validate our approach in a forecasting study for the German day-ahead electricity prices, a highly volatile data set, where we see that the online GAMLSS deliver competitive forecasting accuracy in terms of the CRPS while reducing the computation time by some 1-2 orders of magnitude.
@@
Our research opens up multiple avenues for future research. First, while our empirical results are promising, theoretical results on the regret bounds of the online algorithm would further increase the trust in the proposed method. Secondly, we present a regularized method for the online estimation, but the (online) model selection in distributional regression is a relatively untapped field, and advances here will directly benefit practical applications. Furthermore, our implementation is constrained to the linear parametric GAMLSS so far. While simple splines can be implemented directly in the design matrix, an interesting practical extension would be the ability to incrementally fit penalized splines and automatically select grid points.
CRediT authorship contribution statement
SH: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Writing - original draft, Writing - review & editing, Visualization.
JB: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Writing - original draft, Writing - review & editing, Visualization.
FZ: Conceptualization; Methodology; Validation; Formal analysis; Writing – review and editing; Supervision.
Acknowledgements
Simon Hirsch is employed by Statkraft Trading GmbH and gratefully acknowledges the support and funding received. Simon Hirsch is grateful to Daniel Gruhlke for many helpful discussions. This work contains the author’s opinions and does not necessarily reflect Statkraft’s position. The authors declare no conflict of interest.
Appendix A Relationship between GAMLSS and IRLS
The following section sheds light on the relationship between GAMLSS and the IRLS for fitting regression models for (conditional) heteroskedasticity (see e.g. Dette & Wagener, 2013; Ziel, 2016; Ziel et al., 2016).
For the GAMLSS part, we assume the response to be distributed following the mean-variance parameterization of the normal distribution. This corresponds to the NO2 implementation in the GAMLSS R package. Furthermore, the link functions $g_{1}(\cdot)$ and $g_{2}(\cdot)$ are the identity functions and hence $\eta_{1}=g_{1}(\theta_{1})=\theta_{1}$ and $\eta_{2}=g_{2}(\theta_{2})=\theta_{2}$ For the distribution parameters $\theta_{1}=\mu$ and $\theta_{2}=\sigma^{2}$, derivatives of the likelihood are given by:
@@ Appendix A Relationship between GAMLSS and IRLS
$\dfrac{\partial^{2}l}{\partial\mu^{2}}$ $=-\dfrac{1}{\sigma^{4}}$ $\dfrac{\partial^{2}l}{\partial(\sigma^{2})^{2}}$ $=-\dfrac{1}{2}\dfrac{1}{\sigma^{4}}$
we therefore have the weights $\boldsymbol{W}_{11}=1/\sigma^{2}$ and $\boldsymbol{W}_{11}=1/(2\sigma^{4})$ since $\partial\eta_{k}/\partial\theta_{k}=1$ for the identity link function. The working vector is given by Equation 6 and can be written as:
$z_{1}$ $=z_{\mu}=\mu+\dfrac{\dfrac{1}{\sigma^{2}}(y-\mu)}{\dfrac{1}{\sigma^{2}}}=y$
$z_{2}$ $=z_{\sigma}^{2}=\sigma^{2}+\frac{1}{2}\dfrac{\dfrac{(y-\mu)^{2}-\sigma^{2}}{\sigma^{4}}}{-\left(-\dfrac{1}{2\sigma^{4}}\right)}=(y-\mu)^{2}$
Therefore, the working vectors reduce to the response variable $y$ and the squared residuals $(y-\mu)^{2}$, a natural estimator for the variance. The regression for the mean is weighted by the inverse of the estimated variance.

### v2

@@ 7 Discussion and Conclusion
We validate our approach in a forecasting study for the German day-ahead electricity prices, a highly volatile data set. The online GAMLSS models deliver competitive forecasting accuracy in the CRPS compared to repeated batch estimation and other distributional forecasting approaches while reducing the computation time by some 2-3 orders of magnitude. Furthermore, we provide simulation results on the impact of the initial training size and the exponential discounting on the approximation quality of the online models compared to repeated batch fitting. Our simulation results are in line with the aforementioned theoretical results.
@@
Our research opens up multiple avenues for future research. First, while our empirical results are promising, further theoretical results on the error bounds of the online algorithm would further increase the trust in the proposed method. Secondly, we present a regularized method for the online estimation, but the (online) model selection in distributional regression is a relatively untapped field, and advances here will directly benefit practical applications. Furthermore, our implementation is constrained to the linear parametric GAMLSS so far. While simple splines can be implemented directly in the design matrix, an interesting practical extension would be the ability to fit penalized splines and automatically select grid points incrementally.
Acknowledgments and Disclosure of Funding
Simon Hirsch is employed as an industrial PhD student by Statkraft Trading GmbH and gratefully acknowledges the support and funding received. Simon Hirsch is grateful to Daniel Gruhlke for many helpful discussions. This work contains the author’s opinions and does not necessarily reflect Statkraft’s position. The authors declare no conflict of interest.
Appendix A Proof: Relationship between GAMLSS and IRLS
In the following we prove the result given in Proposition 1. For the parameters $\theta_{1}=\mu$ and $\theta_{2}=\sigma^{2}$, the derivatives of the log-likelihood of the normal distribution are given by:
@@ Appendix A Proof: Relationship between GAMLSS and IRLS
$\dfrac{\partial^{2}l}{\partial\mu^{2}}$ $=-\dfrac{1}{\sigma^{2}}$ $\dfrac{\partial^{2}l}{\partial(\sigma^{2})^{2}}$ $=-\dfrac{1}{2}\dfrac{1}{\sigma^{4}}$
we therefore have the weights $\boldsymbol{W}_{1,1}=1/\sigma^{2}$ and $\boldsymbol{W}_{1,1}=1/(2\sigma^{4})$ since $\partial\eta_{k}/\partial\theta_{k}=1$ for the identity link function. The working vector is given by Equation 6 and can be written as:
$z_{1}$ $=z_{\mu}=\mu+\dfrac{\dfrac{1}{\sigma^{2}}(y-\mu)}{\dfrac{1}{\sigma^{2}}}=y,$ $z_{2}$ $=z_{\sigma^{2}}=\sigma^{2}+\frac{1}{2}\dfrac{\dfrac{(y-\mu)^{2}-\sigma^{2}}{\sigma^{4}}}{-\left(-\dfrac{1}{2\sigma^{4}}\right)}=(y-\mu)^{2}$
which returns the working vectors given in Proposition 1 $\blacksquare$